If an angle is equal to 5times its compliment .Find the measure of the angle.
Answer by LinnW(1048)

Set x = the angle
set c = the complement of x
x + c = 90
Solve for c by adding -x to each side
c = 90 - x
x = 5c
substitute (90 - x) for c
x = 5(90 - x)
x = 450 - 5x
add 5x to each side
6x = 450
divide each side by 6
x = 450/6
x = 75